# Model-Based Transfer Learning for Contextual Reinforcement Learning

### v1

@@ 5.1 Setup
Our experiments span a variety of CMDP settings, focusing on transportation tasks as well as classic control benchmarks. We introduce three intelligent transportation tasks that utilize DRL but suffer from brittleness issues, as well as the classic control tasks.
Baselines. Our proposed methods are compared against a range of baselines including (1) Random selection, where each training task is chosen randomly; (2) Exhaustive training, which involves independent training separate models on all tasks exhaustively but is computationally expensive; (3) Multi-task RL, where a single context-conditioned model is trained for all tasks; (4) Oracle transfer, which assumes full knowledge of generalized performance for all policies and selects the best source task for each target task.
Proposed methods. Our proposed methods are: (1) MBTL-ES, selecting source tasks for the context to be equidistance based on the transfer budget; (2) MBTL-GS, selecting the next source task greedily to maximize the marginal improvement of expected generalized performance based on linear generalization gap (Assumption 3); (3) MBTL-GP, selecting source task based on the BO using GP and UCB with scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$.
@@
DRL algorithms and performance measure. We utilize Deep Q-Networks (DQN) for discrete actions [24] and Proximal Policy Optimization (PPO) for continuous actions [30]. We evaluate our methods by the average performance across all target tasks after training 15 source tasks and the number of trained models needed to achieve a certain level of suboptimality. We employ min-max normalization of the rewards for each task.
5.2 Traffic tasks
Table 1: Comparative performance of different methods on traffic CMDP tasks
Benchmark Baselines MBTL (Ours) Oracle
Task Context Variation Random Exhaustive Multitask MBTL-GS MBTL-ES MBTL-GP Oracle Transfer
Number of Trained Models $k$ $N$ $1$ $k$ $K$ $k$ $N$
Traffic Signal Road Length 0.9249 0.9409 0.8242 0.9278 0.9213 0.9371 0.9409
Traffic Signal Inflow 0.8457 0.8646 0.8319 0.8496 0.8700 0.8673 0.8768
Traffic Signal Speed Limit 0.8821 0.8857 0.6083 0.8862 0.8858 0.8854 0.8876
Eco-Driving Penetration Rate 0.5959 0.5260 0.1945 0.5827 0.5934 0.6323 0.6660
Eco-Driving Inflow 0.4774 0.4061 0.2229 0.4673 0.4705 0.5108 0.5528
Eco-Driving Green Phase 0.4406 0.3850 0.4228 0.4431 0.4557 0.4700 0.5027
AA-Ring-Acc Hold Duration 0.8924 0.8362 0.9209 0.8776 0.9057 0.9242 0.9552
AA-Ring-Vel Hold Duration 0.9785 0.9589 0.9720 0.9807 0.9772 0.9816 0.9822
AA-Ramp-Acc Hold Duration 0.6050 0.4276 0.5158 0.6143 0.5956 0.6318 0.7111
AA-Ramp-Vel Hold Duration 0.6690 0.5473 0.5034 0.5907 0.6787 0.7182 0.7686
Average 0.7312 0.6778 0.6017 0.7220 0.7354 0.7559 0.7844
@@ 5.2 Traffic tasks
†Higher the better. The bold indicates the best for each benchmark, excluding the Oracle transfer. Detailed results with variance for each method are provided in Appendix A.2.2.
@@
We validate our methods on diverse simulated traffic scenarios, focusing on to what extent Our proposed methods can optimize the global objective only with a small number of trained models.First, while most traffic lights operate on fixed schedules, we can design learning-based adaptive (1) Traffic signal control to optimize the traffic [8, 21]. However, considering that every intersection looks different, challenges persist in generalizing across various intersection configurations [18]. Given the significant portion of greenhouse gas emissions in the United States coming from the transportation sector [11], the second traffic task is (2) Eco-driving Lagrangian control, which is critical for climate change mitigation. DRL-based eco-driving strategies have been developed [13, 38, 17] but still have some issues of difficulties in generalization. The last traffic task is (3) Advisory autonomy, which is a real-time speed advisory system that enables human drivers to emulate the system-level performance of autonomous vehicles in mixed autonomy systems [34, 15]. Instead of direct and instantaneous control, human drivers receive periodic guidance, which varies based on the traffic states. In Appendix A.2, we provide comprehensive details about our experiments.
Figure 5: Comparison of normalized generalized performance of all target tasks.
Discussion. Table 1 compares the min-max normalized generalized performance across the full target task range. While exhaustive training is computationally expensive and exhibits suboptimal performance, the Oracle transfer, which assumes perfect knowledge of transferability, demonstrates the potential benefits of zero-shot transfer for solving the complete range of multiple tasks. Our proposed MBTL methods leverage this transferability by intelligently selecting the next source task to learn from. Notably, MBTL-GP shows superior performance, closely aligning with the Oracle transfer, which suggests the Gaussian Process effectively models the underlying task dynamics, facilitating better generalization from learned experiences. For example, MBTL-GP significantly improves generalized performance with a few source tasks for traffic signal control problems varying green phase durations (Figure 5). More results are provided in Appendix A.2.
5.3 Standard control benchmarks
@@
Table 2: Comparative performance of different methods on standard control CMDP tasks
Benchmark Baselines MBTL (Ours) Oracle
Task Context Variation Random Exhaustive Multitask MBTL-GS MBTL-ES MBTL-GP Oracle Transfer
Number of Trained Models $k$ $N$ $1$ $k$ $K$ $k$ $N$
Pendulum Length 0.7270 0.7383 0.6830 0.7327 0.7092 0.7697 0.7969
Pendulum Mass 0.6329 0.6237 0.5793 0.6408 0.6092 0.6827 0.7132
Pendulum Timestep 0.7989 0.8135 0.7247 0.8177 0.7541 0.8141 0.8801
Cartpole Mass of Cart 0.7221 0.9466 0.7153 0.6501 0.7516 0.8212 0.9838
Cartpole Length of Pole 0.8121 0.9110 0.5441 0.8217 0.8428 0.9124 0.9875
Cartpole Mass of Pole 0.8858 0.9560 0.6073 0.8744 0.7909 0.9351 1.0000
BipedalWalker Gravity 0.9330 0.9281 0.7898 0.9359 0.9494 0.9393 0.9674
BipedalWalker Friction 0.9650 0.9317 0.9051 0.9645 0.9664 0.9713 0.9778
BipedalWalker Scale 0.8605 0.8694 0.7452 0.8792 0.8496 0.8886 0.9107
HalfCheetah Gravity 0.8542 0.6679 0.6292 0.8663 0.8634 0.9073 0.9544
HalfCheetah Friction 0.8567 0.6693 0.7242 0.8591 0.8703 0.9274 0.9663
HalfCheetah Stiffness 0.8533 0.6561 0.7007 0.8785 0.7817 0.9146 0.9674
Average 0.8251 0.8093 0.6957 0.8267 0.8116 0.8736 0.9255
@@
†Higher the better. The bold indicates the best for each benchmark, excluding the Oracle transfer. Detailed results with variance for each method are provided in Appendix A.2.2.
@@
Building on the evaluation of a suite of traffic-related tasks, we extend our investigation to classic control tasks to assess the applicability of our methods across a broader spectrum of environments. We utilize context-extended versions of standard RL environments from CARL benchmark library [6] to rigorously test the robustness and adaptability of our methods under varied contexts. For Cartpole tasks, we explored contextual Markov decision processes (CMDPs) with varying cart masses, pole lengths, and pole masses. We vary the timestep duration, pendulum length, and pendulum mass in Pendulum. The BipedalWalker was tested under varying friction, gravity, and scale. In HalfCheetah tasks, we manipulated friction, gravity, and stiffness parameters. These variations critically influence the dynamics and physics of the environments. The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times, enabling a comprehensive analysis of transfer learning under drastically different conditions. We provide more details in Appendix A.2.
Discussion. Table 2 compares the performance across the full target task range. Our experiments demonstrate that MBTL significantly improves performance and sample efficiency in dynamic urban transportation systems compared to the baselines. In control tasks such as Cartpole, Pendulum, BipedalWalker, and HalfCheetah, where physical parameters like mass directly influence the dynamics, MBTL methods, particularly MBTL-GP, excel by effectively transferring knowledge between similar task variations. For example, in Figure 5, MBTL-GP outperforms other baselines, including multitask training. The results underscore the potential of MBTL in solving multiple related tasks.
Figure 6: Sensitivity analysis on different DRL algorithms (Cartpole, length of pole).
Figure 7: Sensitivity analysis on the different acquisition functions.
@@ 5.3.1 Sensitivity analysis
DRL algorithms. Figure 6 shows an ablation study comparing different DRL algorithms—DQN, PPO, and Advantage Actor-Critic (A2C) [25]. The heatmap results show differences in transferability, and the results underscore the robustness of MBTL to different algorithms.
Acquisition functions. Figure 7 assesses the role of acquisition functions in Bayesian optimization. While expected improvement (EI) focuses on promising marginal gains beyond the current best, UCB utilizes both mean and variance for balancing exploration and exploitation. For Cartpole and advisory autonomy tasks, UCB with decreasing $\beta_{k}$ as $k$ increases shows the best performance. However, for eco-driving control, the $\beta_{k}$ formulation in Section 4.4 exhibits the highest performance, though the difference is not statistically significant.
@@ 6 Related work
Robustness and generalization challenges in DRL are generally addressed by a few common techniques in the literature. The broader umbrella of such methods falls under CRL [6]. CRL utilizes side information about the problem variations to improve the generalization and robustness. In particular, CRL formalizes generalization in DRL using CMDPs [14, 27, 6], which incorporate context-dependent dynamics, rewards, and initial state distributions into the formalism of MDPs. This provides a structured framework to study generalization in DRL [14, 27].
The contexts of CMDPs are not always visible during training [20]. When they are visible, they can be directly used as side information by conditioning the policy on them [33]. A common approach is multi-task learning, which aims to leverage a common structure between tasks. The structure between tasks has been exploited using policy sketches for task structuring [2] and by sharing a distilled policy that captures common behavior across tasks [36]. However, when the context features are not visible, the resulting CMDP becomes a partially observable CMDP [20, 9] and can be challenging for multi-task learning. In this work, we use multi-task learning as a baseline in evaluating our methods.
Zero-shot transfer is another commonly used technique that focuses on adapting models trained in one environment to perform effectively in unseen settings without further training [20]. Rezaei-Shoshtari et al. [29] utilize hypernetworks to approximate an RL algorithm as a mapping from parameterized CMDP to a family of near-optimal solutions to achieve zero-shot transfer across varying task conditions. Sinapov et al. [32] uses meta-data to learn inter-task transferability to learn the expected benefit of transfer given a source-target task pair.
Another popular class of techniques for robust and generalizable DRL is domain adaptation strategies. It is often used in making learned policies adapt to new domains with varying dynamics, observations, or rewards [5]. For instance, Xing et al. [39] first learn a latent unified state representation for different domains and subsequently train DRL agents in the source domain based on that. However, our focus in this work centered around in-distribution generalization, whereas domain adaptation techniques are commonly used for out-of-distribution generalization.
@@ 7 Conclusion
Our study refines the robustness of DRL by developing MBTL using zero-shot transfer. MBTL strategically selects a set of source tasks to maximize overall performance and minimize training costs. It models the generalization gap and uses BO to predict task performance efficiently. We theoretically show that the method exhibits regret that is sublinear in the number of training tasks. We validate MBTL with urban traffic and control benchmarks. Limitation is that MBTL relies highly on assumptions that may not hold in highly dynamic or complex environments, potentially limiting its applicability and leading to suboptimal policy convergence. Future work will focus on improving the modeling of the generalization gap and exploring scenarios with multiple context variations where simple strategies like MBTL-ES or MBTL-GS may not be applicable.

### v2

@@ 5.1 Setup
Our experiments consider CMDPs that span standard and real-world benchmarks. In particular, we consider standard continuous control benchmarks from the CARL library [6]. In addition, we study problems from RL for intelligent transportation systems, using [47] to model the CMDPs. Surprisingly, despite the relatively low complexity of the CMDPs considered, standard deep RL algorithms appear to struggle to solve the tasks.
Baselines. We consider two types of baselines when evaluating our proposed algorithm: canonical and multi-policy. The canonical baselines are selected to validate multi-policy training; the multi-policy training baselines are heuristic methods designed to validate the Bayesian optimization approach. The canonical baselines include: (1) Independent training, which involves independently training separate models on each task; and (2) Multi-task RL, where a single “universal” context-conditioned policy is trained for all tasks. The multi-policy baselines include: (3) Random selection, where each training task is chosen uniformly at randomly; (4) Equidistant strategy, which selects training tasks by equally subdividing the context space based on a given training budget, and then trains them in lexicographical order; (5) Greedy strategy, which greedily selects the next source task based on Assumption 1 and fixed training performance; and (6) Sequential Oracle transfer, which has access to generalized performance corresponding to policies for all tasks (including those not yet selected) and uses that information to greedily select the best source task at each step.
Proposed method. We evaluate MBTL with the scaling factor $\beta_{k}=2\log(|X|\pi^{2}k^{2}/6\delta)$.
@@
DRL algorithms and performance measure. We utilize Deep Q-Networks (DQN) for discrete action spaces [29] and Proximal Policy Optimization (PPO) for continuous action spaces [35]. We evaluate our method by the average performance across all $N$ target tasks after training up to $K=$15 source tasks or the number of source tasks needed to achieve a certain level of performance. We employ min-max normalization of the rewards for each task, and we provide comprehensive details about our model in Appendix A.4.1.
5.2 Traffic benchmark experiments
Table 1: Comparative performance of different methods on traffic CMDPs
Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
Traffic Signal Inflow 0.8646 0.8319 0.8457 0.8700 0.8496 0.8673 0.8768
Traffic Signal Speed Limit 0.8857 0.6083 0.8821 0.8858 0.8862 0.8854 0.8876
Eco-driving Penetration Rate 0.5260 0.1945 0.5959 0.5934 0.5827 0.6323 0.6660
Eco-driving Inflow 0.4061 0.2229 0.4774 0.4705 0.4673 0.5108 0.5528
Eco-driving Green Phase 0.3850 0.4228 0.4406 0.4557 0.4431 0.4700 0.5027
AA-Ring-Acc Hold Duration 0.8362 0.9209 0.8924 0.9057 0.8776 0.9242 0.9552
AA-Ring-Vel Hold Duration 0.9589 0.972 0.9785 0.9772 0.9807 0.9816 0.9822
AA-Ramp-Acc Hold Duration 0.4276 0.5158 0.6050 0.5956 0.6143 0.6318 0.7111
AA-Ramp-Vel Hold Duration 0.5473 0.5034 0.669 0.6787 0.5907 0.7182 0.7686
Average 0.6778 0.6017 0.7312 0.7354 0.722 0.7559 0.7844
@@ 5.2 Traffic benchmark experiments
†Higher the better. Bold values represent the highest value(s) within the statistically significant range for each CMDP, excluding the oracle. Detailed results with variance for each method are provided in Appendix A.4.2.
‡AA: Advisory autonomy benchmark, Ring: Single lane ring, Ramp: Highway ramp, Acc: Acceleration guidance, Vel: Speed guidance.
@@
We consider three traffic control domains. First, while most traffic lights still operate on fixed schedules, RL can be used to design adaptive (1) Traffic signal control to optimize traffic [9, 23]. However, considering that every intersection is different, challenges persist in generalizing across intersection configurations [19]. Given the significant portion of greenhouse gas emissions in the United States due to transportation [12], the second traffic domain is (2) Dynamic eco-driving at signalized intersections, which concerns learning energy-efficient driving strategies in urban settings. DRL-based eco-driving strategies have been developed [14, 45, 18] but still experience difficulties in generalization. Our final traffic domain is (3) Advisory autonomy, in which real-time speed or acceleration advisories guide human drivers to act as vehicle-based traffic controllers in mixed traffic environments [39, 8, 16]. The context space $X$ is discretized into $N=\{50,50,40\}$ contexts for the three domains, respectively. In Appendix A.4, we provide details about our experiments.
Figure 4: Traffic CMDP results. Method comparison of normalized performance over $N$ tasks. MBTL efficiently selects source training tasks. The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 25x fewer samples needed.
Results. Table 1 and Figure 4 summarize the results. Notably, the Oracle far outperforms the standard baselines (Independent and Multi-task), indicating the potential for transfer learning and intelligent training of multiple models, respectively. MBTL rapidly approaches the Oracle, within $\approx 10$ transfer steps, indicating that the Gaussian process effectively models the training performance and linear generalization gap models generalization performance. It is important to note that multi-task RL studies commonly consider Independent training as a strong baseline due to its avoidance of negative transfer and other training instability issues. Indeed, Independent often (but not always) outperforms Multi-task in our experiments. Yet, our experiments show that MBTL consistently outperforms both Independent and Multi-task and matches their performance with 5-25x better sample efficiency. Among the sequential baselines, MBTL consistently outperforms the heuristic baselines, indicating the value of adaptively selecting source tasks based on feedback. The sequential baselines, such as random, equidistant, and greedy strategy, also generally outperform Independent and Multi-task, indicating the general value of multi-policy training for solving CMDPs. More results are provided in Appendix A.4.
5.3 Continuous control benchmark experiments
@@
Table 2: Comparative performance of different methods on standard control CMDPs
Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
Pendulum Length 0.7383 0.6830 0.7270 0.7092 0.7327 0.7697 0.7969
Pendulum Mass 0.6237 0.5793 0.6329 0.6092 0.6408 0.6827 0.7132
Pendulum Timestep 0.8135 0.7247 0.7989 0.7541 0.8177 0.8141 0.8801
Cartpole Mass of Cart 0.9466 0.7153 0.7221 0.7516 0.6501 0.8212 0.9838
Cartpole Length of Pole 0.9110 0.5441 0.8121 0.8428 0.8217 0.9124 0.9875
Cartpole Mass of Pole 0.9560 0.6073 0.8858 0.7909 0.8744 0.9351 1.0000
BipedalWalker Gravity 0.9281 0.7898 0.9330 0.9494 0.9359 0.9393 0.9674
BipedalWalker Friction 0.9317 0.9051 0.965 0.9664 0.9645 0.9713 0.9778
BipedalWalker Scale 0.8694 0.7452 0.8605 0.8496 0.8792 0.8886 0.9107
HalfCheetah Gravity 0.6679 0.6292 0.8542 0.8634 0.8663 0.9073 0.9544
HalfCheetah Friction 0.6693 0.7242 0.8567 0.8703 0.8591 0.9274 0.9663
HalfCheetah Stiffness 0.6561 0.7007 0.8533 0.7817 0.8785 0.9146 0.9674
Average 0.8093 0.6957 0.8251 0.8116 0.8267 0.8736 0.9255
@@
†Higher the better. Bold values represent the highest value(s) within the statistically significant range for each CMDP, excluding the oracle. Detailed results with variance for each method are provided in Appendix A.4.2.
Figure 5: Continuous control CMDP results. Method comparison of normalized performance over $N$ tasks. The black dotted line indicates the first training step within MBTL that exceeds both independent and multi-task baselines, with up to 50x fewer samples needed.
@@
To probe the generality of MBTL, we utilize context-extended versions of standard RL environments from CARL benchmark library [6] to evaluate our methods under varied contexts. For the Cartpole domain, we considered CMDPs with varying cart mass, pole length, and pole mass. In Pendulum, we vary the timestep duration, pendulum length, and pendulum mass. The BipedalWalker was tested under varying friction, gravity, and scale. In HalfCheetah domain, we manipulated friction, gravity, and stiffness parameters. These variations critically influence the dynamics and physics of the environments. The range of context variations was selected by scaling the default values specified in CARL from 0.1 to 10 times ($N=100$), enabling a comprehensive analysis of transfer learning under drastically different conditions. We provide more details in Appendix A.4.
Figure 6: The GP sequentially updates estimates of the performance function (blue) based on previously trained models. Then, MBTL selects the next source task that maximizes the acquisition function (red). (CMDP: Pendulum (Time step)).
Results. The results summarized in Table 2 demonstrate MBTL’s sample efficiency and competitive performance across diverse control domains, often closely trailing the Oracle only with a small number of tasks. With the exception of the Cartpole domain, MBTL generally shows superior performance. Figure 5 shows that MBTL demonstrates rapid learning curves, reaching high normalized performance with relatively few samples. Specifically, Figure 6 illustrates the detailed process of how MBTL utilizes GP for performance estimation and chooses the next source task that maximizes the acquisition function that evaluates the expected improvement of generalized performance. MBTL achieves comparable performance to multi-task or independent baselines with 8-50x fewer samples, highlighting its effectiveness in reducing training requirements.
Figure 7: Sensitivity analysis on the DRL algorithm underlying MBTL (DQN, PPO, and A2C), tested on Cartpole with varying length of pole. MBTL remains effective.
Figure 8: Sensitivity analysis on acquisition functions.
@@ 5.3.1 Sensitivity analysis
DRL algorithms. Figure 7 shows that MBTL remains effective with different underlying DRL algorithms—DQN, PPO, and Advantage Actor-Critic (A2C) [30]—used for single-task training.
Acquisition functions. Figure 8 assesses the role of acquisition functions in Bayesian optimization. While expected improvement (EI) focuses on promising marginal gains beyond the current best, UCB utilizes both mean and variance for balancing exploration and exploitation. Overall, we find that MBTL is not particularly sensitive to the choice of optimism representation in the acquisition function, which indicates that MBTL has a weak dependence on hyperparameters.
@@ 6 Related work
Contextual Reinforcement Learning. Robustness and generalization challenges in DRL are generally addressed by a few common techniques in the literature. The broader umbrella of such methods falls under CRL [6]. CRL utilizes side information about the problem variations to improve the generalization and robustness. In particular, CRL formalizes generalization in DRL using CMDPs [15, 32, 6], which incorporate context-dependent dynamics, rewards, and initial state distributions into the formalism of MDPs. This provides a structured framework to study generalization in DRL [15, 32]. The contexts of CMDPs are not always visible during training [22]. When they are visible, they can be directly used as side information by conditioning the policy on them [38].
Multi-task learning. A common approach for solving CRL is multi-task learning, which aims to leverage common structure between tasks. The structure between tasks has been exploited using policy sketches for task structuring [2] and by sharing a distilled policy that captures common behavior across tasks [44]. One of the known challenges in multi-task learning is that when the context features are not visible, the resulting CMDP becomes a partially observable CMDP [22, 10] and can be challenging for multi-task learning. Moreover, another challenge is negative transfer, where training from tasks that are too different results in training instability or failure [21, 41, 42]. In this work, we use multi-task learning as a baseline in evaluating our methods.
Zero-shot transfer and policy reuse. Zero-shot transfer is another commonly used technique that focuses on adapting models trained in one environment to perform effectively in unseen settings without further training [22]. Li et al. [25] introduces context-aware policy reuse, which adapts multiple source policies based on context, enhancing generalization in complex environments. Rezaei-Shoshtari et al. [34] utilize hypernetworks to approximate an RL algorithm as a mapping from parameterized CMDP to a family of near-optimal solutions to achieve zero-shot transfer across varying task conditions. Sinapov et al. [37] use meta-data to learn inter-task transferability to learn the expected benefit of transfer given a source-target task pair. Bao et al. [3] propose a metric for evaluating transferability based on information-theoretic feature representations across tasks. Another popular class of techniques for robust and generalizable DRL is domain adaptation strategies. It is often used in making learned policies adapt to new domains with varying dynamics, observations, or rewards [5]. For instance, Xing et al. [46] first learn a latent unified state representation for different domains and subsequently train DRL agents in the source domain based on that. However, our focus in this work centered around in-distribution generalization, whereas domain adaptation techniques are commonly used for out-of-distribution generalization.
Source task selection. In the context of transfer learning, selecting appropriate source tasks is crucial. Li and Zhang [24] proposes an optimal online method for dynamically selecting the most relevant single source policy in reinforcement learning. Beyond RL, Meiseles and Rokach [28] emphasizes structural alignment in time-series source models to prevent performance degradation, while Poth et al. [33] finds that selecting aligned intermediate tasks in natural language processing boosts transfer effectiveness. Together, these studies show the importance of context-aligned source tasks selection to improve transferability and reduce negative transfer.
@@ 7 Conclusion
This study introduces a method called Model-based Transfer Learning (MBTL), which layers on top of existing RL methods to effectively solve CMDPs. Rather than independent or multi-task training, which trains $N$ or 1 models, respectively, MBTL intelligently selects an intermediate number of models to train. MBTL has two key components: an explicit model of the generalization gap and a Gaussian process component to estimate training performance. MBTL achieves up to 50x improved sample efficiency on standard and real-world benchmarks. Furthermore, MBTL achieves sublinear regret in the number of training tasks. A limitation is that MBTL is designed for a single-dimensional context variation. Promising directions of future work include studying high-dimensional context spaces and out-of-distribution generalization, as well as the development of new real-world CMDP benchmarks.
Acknowledgments and Disclosure of Funding
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported within this paper. This work was supported by the National Science Foundation (NSF) CAREER award (#2239566), the Kwanjeong Educational Foundation Ph.D. scholarship program, and an Amazon Robotics Ph.D. Fellowship. The authors would like to thank the anonymous reviewers for their valuable feedback.